OpenStudy (anonymous):

Bart bought a digital camera with a list price of $219 from an online store offering a 6 percent discount. He needs to pay $7.50 for shipping. What was Bart's total cost?

OpenStudy (anonymous):

\[219*\frac{94}{100}+\frac{750}{100}\text{=}\frac{5334}{25}=\text{$\$$213}.36\]